OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) A y equals negative two-thirds x plus 3 B y equals negative two-thirds x plus five-thirds C y = 2x – 21 D y equals negative two-thirds x minus 9

OpenStudy (mathstudent55):

1. Solve the given equation for y. You'll get an equation of the form y = mx + b. m is the slope. 2. The line that is parallel to it has the same slope, m. To get its equation, use the form of the equation y - y1 = m(x - x1), where (x1, y1) is the known point and m is the slope you obtained above. Substitute x1, y1, and m into this equation to get the equation of the parallel line.